OpenStudy (sonyalee77):

Solve the linear system by graphing. y=-x-2 y=x+6

OpenStudy (anonymous):

if you'll graph these two functions, there will be an intersection. This intersection will be the solution of this system.

OpenStudy (sonyalee77):

but how?

OpenStudy (sonyalee77):

@wolfe8

OpenStudy (anonymous):

For the first one, you'll go right 2 but make it facing down since it's a negative x. For the second, you go left 6.

OpenStudy (yanasidlinskiy):

Set them up equal to each other like this: -x-2=x+6 -------To solve it you have to get one variable to the other side. either one +x +x -2 = x+6 --------To get x by itself, always do the opposite of what is With the number -6 -6 - and + cancel each other out now you are left with x -8 = x ------> divide is your last step And your answer is x=-8

OpenStudy (sonyalee77):

but how do you get the y coordinate

OpenStudy (yanasidlinskiy):

Here is how: You have your x. Now to get your y, all you have to do is plug in x for the equation. Doesn't matter which one:) y=x+6 y=-8+6 y=-2

OpenStudy (sonyalee77):

so it is -8,-2

OpenStudy (yanasidlinskiy):

Yep!!:) You're correct!!:) Do you see where I got that from?